4. Place parentheses, if needed, to make each of the following equations true:

a. 4 + 3 x 2 = 14
b. 9 ÷ 3 + 1 = 4
c. 5 + 4 + 9 ÷ 3 = 6
d. 3 + 6 - 2 ÷ 1 = 7

Found 2 solutions by richard1234, karaoz:
Answer by richard1234(7193)

(4 + 3) x 2 = 14

(9 / 3) + 1 = 4

(5 + 4 + 9) / 3 = 6

(3 + 6 - 2) / 1 = 7

Answer by karaoz(32)


Correct solution to your problem is:

a. (4 + 3) × 2 = 14
Explanation: The original expression is 4 + 3 × 2, which is equal to 4 + 6 = 10, which is not equal to 14. Hence we need to change something to make it equal to 14. There is only one possible change in the order of operations since there were only two operations in total. Hence (4 + 3) × 2 = 14. And it worked.

b. 9 ÷ 3 + 1 = 4
Explanation: The given equation is true and therefore no parentheses should be placed.

c. (5 + 4 + 9) ÷ 3 = 6
Explanation: The original equation boils down to 12 = 6, which is not true so some change in the order of operations was necessary. Placing parentheses around 5 + 4 + 9 make the sum equal to 18, which when divided by 3 gives exactly 6.

d. 3 + 6 - 2 ÷ 1 = 7
Explanation: The given equation is true and therefore no parentheses should be placed.